The population mean weight of a candy is 10 grams with a standard deviation of 2 grams. A sample of 100 candies is taken. What is the probability that the sample mean differs from the population mean by at most 0.3 gram?
1
Expert's answer
2020-10-26T19:36:30-0400

Given:

"\mu=10\\\sigma=2\\n=100\\"

ind the probability that the mean of a sample of size 100 will differ from the population mean 10 by at most 0.3 unit, that is, is either more than 9.7 or less than 10.3The corresponding z-value needed to be computed

Formula: "Z= \frac{\overline{X}-\mu}{\frac{\sigma }{\sqrt{n}}}" = "\frac{9.7-10}{\frac{2 }{\sqrt{100}}}" = -1.5

Z= "\frac{10.3-10}{\frac{2 }{\sqrt{100}}}" =1.5


Using z table,"Pr(9.7≤ 
\overline{X}
 ≤10.3)" "=Pr(\u22121.5\u2264Z\u22641.5)"

"=Pr(Z\u22641.5)\u2212Pr(Z\u2264\u22121.5)"

=0.9332−0.0668=0.8664